A company that manufactures large LCD screens knows that not all pixels on their screen​ light, even if they spend great care when making them. They know that the manufacturing process has a blank pixel​ rate, on​ average, of 4.4 blank pixels in a sheet 4 ft by 12 ft that will be cut into smaller screens. They believe that the occurrences of blank pixels are independent. Their warranty policy states that they will replace any screen sold that shows more than 2 blank pixels. Suppose the number of blank pixels can be modeled by a Poisson distribution. Complete parts​ a) through​ d) below.

​c) What is the probability that a 2 ft by 3 ft screen will have at least one​ defect?

A 2 ft by 3 ft screen will have at least one defect with probability _____.

​(Round to three decimal places as​ needed.)
Answer #1

c)

expected number of defect on 2 ft by 3 ft screen =4.4*(2*3)/(4*12)=0.55

P( A 2 ft by 3 ft screen will have at least one defect with probability )=P(X>=1)=1-P(X=0)

=1-e-0.55 =0.423